OpenStudy (anonymous):

What is 4x + 2y = -20 written in slope-intercept form? A. y = 2x - 20 B. y = -4x + 20 C. y = 20x - 4 D. y = -2x - 10

OpenStudy (anonymous):

is d , slope intercept form equals y= mx+b, so y=4x+2y=-20, solve for y, meaning leave y in one side and all the other variable in the other side. subtract 4x, so ur lefts with 2y= -20 -4x, then divide both sodes by 2 to leave y alone. so y=-20/2 and -4x/2 which equals to y= -10-2x, now organized it to slope from, y= mx+b becomes y=-2x-10.